# Symplectic torus actions on complete intersections

### v1

@@ Abstract
In [20] the author classified complete intersections of dimension $8k$ admitting a Hamiltonian circle action, under a certain assumption on the fixed point set. In this paper, we improve this result in two directions. Firstly, in dimension $8$ we remove the assumption on the fixed point set. Secondly, in any dimension we prove the result under an analogous assumption on the fixed point set. We also give some applications of our methods to the unimodality of Betti numbers of symplectic manifolds having a Hamiltonian $S^{1}$-action, and discuss the relation to symplectic rationality problems.
@@ 1 Introduction
To state our results, we set the following notation for complete intersections. Let $X_{m}(d_{1},\ldots,d_{k})$ denote a smooth complete intersection in $\mathbb{CP}^{k+m}$ with multidegree $(d_{1},\ldots,d_{k})$. Here, $m$ is equal to the complex dimension, and $k$ the complex codimension.
@@
Since complete intersections of dimension at least $4$ are simply connected, there is no distinction between Hamiltonian and symplectic circle actions. To shorten the statement of results, we keep the understanding that a complete intersection $(M,\omega)$ is a closed symplectic manifold diffeomorphic to a complete intersection. For some results appearing later in the article, some more restrictions on the symplectic form are needed and these are stated explicitly.
@@
The major motivating result of this paper is a theorem of Benoist which gave a complete characterization of smooth complete intersections whose holomorphic automorphism groups are infinite [1, Theorem 3.1]. The aim of this work is to obtain a symplectic analogue of this theorem. A comprehensive smooth analogue in dimension $6$ was given by Dessai and Wiemeler [8].
@@
The first main result is to remove the technical assumption on the fixed point set in our previous result on complete intersections [20, Theorem 1.1] in dimension 8.
@@
Suppose that $(M,\omega)$ is an $8$-dimensional complete intersection with a symplectic circle action. Then, $M$ is diffeomorphic to either $X_{4}(1),X_{4}(2)$ or $X_{4}(2,2)$.
The main issue to overcome in the proof is to deal with $4$-dimensional fixed components. For this we apply a result of $4$-dimensional symplectic geometry due to Ohta and Ono (building on fundamental work of Gromov, McDuff, and Taubes), to deduce information about their intersection form, and deduce via localisation formulas, information about the intersection form of $M$.
In the next part of the paper, we prove a classification result for symplectic circle actions on complete intersections under a condition on the fixed point set analogous to [20, Theorem 1.1], therefore completely removing the assumption on the dimension.
@@
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic circle action with no fixed components $F$ satisfying $\dim(F)>0$ and $\dim(M)=\dim(F)\mod 4$.
Then $M$ is diffeomorphic to $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ or $X_{2}(3),$ $m\in\mathbb{N}$.
Theorem 1.2 has the following corollary.
@@
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic circle action having only isolated fixed points. Then $M$ is diffeomorphic to $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ or $X_{2}(3),$ $m\in\mathbb{N}$.
Projective spaces and quadrics of any dimension admit a Hamiltonian torus action with isolated fixed points. The underlying smooth $4$-manifolds of the complete intersections $X_{2}(2,2),$ or $X_{2}(3),$ admit a symplectic form with a toric action, since both are diffeomorphic to the blow-up of $\mathbb{CP}^{2}$ in a sequence of points.
Hence, the only unknown case remaining from Theorem 1.3 is $X_{2m}(2,2)$ for $m\geq 2$. We are able to deal with this when the action is GKM, and the symplectic form is homotopic to the Fubini-Study form, which we denote $\omega_{FS}$. Thus obtaining in Theorem 4.2, a Benoist type “if and only if” classification of complete intersections having such symplectic torus actions, precisely when the holomorphic automorphism group, denoted $Aut(M)$ is infinite. The proof uses results from [14], to evaluate the integral of $c_{1}$ along a hypothetical toric $1$-skeleton Poincaré dual to $c_{n-1}$, and thus obtaining a contradiction with the known Chern numbers of the intersection of two quadrics.
@@
In the final section of the paper, we prove some new results regarding the unimodality of Betti numbers of a closed symplectic manifold admitting a Hamiltonian circle action. Tolman has asked whether for a closed symplectic manifold with a Hamiltonian circle action, the even Betti numbers are unimodal (i.e. weakly increasing up to the middle degree) [15, Problem 4.3]. For Kähler manifolds the property follows from the hard Lefschetz theorem. [15, Problem 4.3] also asked about the sequence of odd Betti numbers, but the results of this paper are only relevant for the even case.
In [5] Cho and Kim gave an affirmative answer to [15, Problem 4.3 (1)] when $M$ is $8$-dimensional and the action has isolated fixed points. In Theorem 5.1 we prove the unimodality of even Betti numbers for $8$-dimensional manifolds having a Hamiltonian $S^{1}$-action with no extremal, 4-dimensional fixed components.
@@
Suppose that $M$ is a closed symplectic $8$-manifold having a Hamiltonian $S^{1}$-action, and neither $M_{\min}$ or $M_{\max}$ are $4$-dimensional. Then $b_{2}(M)\leq b_{4}(M)$.
In Theorem 5.1 we prove that for a closed symplectic $12$-manifold with $b_{2}(M)=1$, satisfying a similar assumption on the fixed point set the even Betti numbers are unimodal. To the author‘s knowledge this is new under the assumption of isolated fixed points.
Furthermore, we give versions of the above results in general dimensions, generalizing an inequality of Cho [3] on the Betti numbers of a manifold admitting a Hamiltonian $S^{1}$-action with isolated fixed points , to more general fixed point sets. Finally, we combine this result with the Rokhlin-Ochanine theorem to obtain a final application.
Acknowledgments. I would like to thank Leonor Godinho and Silvia Sabatini for helpful discussions, in particular explaining to me the results of [14], which allowed me to deal with the intersection of two quadrics in the GKM case. I presented some of the results about unimodality of Betti numbers in the workshop “Perspectives in equivariant topology” in Cologne September 2022, I would like the thank those present for interesting discussions, in particular Liat Kessler and Susan Tolman. I would like to thank Dmitri Panov for collaborations and discussions which influenced the overall approach of the project, and for comments on a draft version.
@@ 2 Preliminaries
2.1 Hamiltonian circle actions
Firstly, we recall the definition of Hamiltonian torus action on a symplectic manifold. Let $\mathfrak{t}$ denote the Lie algebra of the torus $T^{k}=T$, and $\mathfrak{t}^{*}$ its dual. Recall that if a manifold $M$ has a smooth $T$-action, then for every $t\in\mathfrak{t}$ we can naturally associate an invariant vector field on $M$; $X_{t}$, tangent to the orbits of the action of the one-parameter family associated to $t$.
@@ 2.1 Hamiltonian circle actions
Let $(M,\omega)$ be a closed symplectic manifold, with a smooth $T^{k}$-action by symplectomorphims. Then, the action is called Hamiltonian if there is a smooth function $\mu:M\rightarrow\mathfrak{t}^{*}$ such that
@@
$\mu$ is called the moment map of the action. The fixed point set of the action, denoted $M^{T}$ is equal to the critical set of $\mu$, and is a collection of symplectic submanifolds.
When $k=1$ the moment map is ussually referred to as the Hamiltonian and denoted $H:M\rightarrow\mathbb{R}$. The Hamiltonian $H$ is a perfect Morse-Bott function [17], and a perfect Morse function if and only if $M^{S^{1}}$ is finite.
Let $G\subset S^{1}$ be a subgroup. The subset $M^{G}=\{x\in M|\;g.p=p\;\forall g\in G\}$ is an invariant symplectic submanifold for which the action is Hamiltonian, with Hamiltonian $H|_{M^{G}}$. This will be used in the proof of Theorem 1.1. Connected components of $M^{G}$ are called isotropy submanifolds.
For a fixed component $F$ denote by $\lambda_{F}$ half the Morse-Bott index of $H$ along $F$. It holds that $\lambda_{F}$ is also equal to the the number of strictly negative weights of the $S^{1}$-action along $F$ counted with multiplicity. We will need the following localisation formula for the Betti numbers.
@@
[17] Let $(M,\omega)$ be a closed symplectic manifold having a Hamiltonian circle action. Then for each $i$,
$b_{i}(M)=\sum_{F\subset M^{S^{1}}}b_{i-2\lambda_{F}(M)}(F),$
where the sum runs over connected components of the fixed point set.
The following localisation formula is a consequence of the equivariant version of the Atiyah-Singer index theorem. The formula we will use is derived in [12, Section 5.8]. We use $\sigma(M)$ to denote the signature of $M$, which is equal to zero for manifolds of dimension not divisible by $4$ and equal to $1$ for a single point.

### v2

# Cohomological localization for Hamiltonian $S^{1}$-actions and symmetries of complete intersections

@@ Abstract
To begin the paper we revisit a cohomological localization result of Jones-Rawnsley which was subsequently improved by Farber, further generalizing the result. We then proceed to improve a previous result of the author on complete intersections of dimension $8k$ with a Hamiltonian $S^{1}$-action in two directions. Firstly, in dimension $8$ we remove the assumption on the fixed point set. Secondly, in any dimension we prove the result under an analogous assumption on the fixed point set. We also give some applications towards the unimodality of Betti numbers of symplectic manifolds having a Hamiltonian $S^{1}$-action, and discuss the relation to symplectic rationality problems.
@@ 1 Introduction
This paper is concerned with the topology of closed symplectic manifolds having a Hamiltonian $S^{1}$-action. A driving question in symplectic geometry is to understand the extent to which the topology of closed symplectic manifolds differ from that of their Kähler counterparts, of which complex projective varieties are a fundamental source of examples. In general the difference is substantial. For instance, the fundamental group of a closed Kähler manifold satisfies non-trivial restrictions, the most basic being that the abelianization has even rank. On the other hand every finitely presented group appears as the fundamental group of a closed symplectic $4$-manifold [16, Theorem 1.1]. The equivariant version of the problem is an interesting refinement since the presence of symmetries may eliminate the difference between the categories. Key examples of such results are given by Karshon’s classification of closed symplectic $4$-manifolds with a Hamiltonian $S^{1}$-action [24], and Tolman’s classification of closed symplectic $6$-manifolds having a Hamiltonian $S^{1}$-action and $b_{2}=1$ [39]. Here classification has a slightly different meaning in each case, we refer the reader to the original papers for the precise results.
We start by revisiting a result by Jones and Rawnsley [23, Theorem 1], subsequently improved by Farber [12, Corollary 6], about the signature and Betti numbers of a closed symplectic manifold having a Hamiltonian $S^{1}$-action. These results fit into the above discussion, because they generalize results that were well-known in the complex projective category to the symplectic category under the assumptions of symmetry. See the footnote at the beginning of Section 3 for a precise statement.
After recalling some preliminaries in Section 2, we begin Section 3 by proving Proposition 3.2, a localization formula for the following topological invariant:
$I_{JR}(X):=\sigma(X)-\sum_{i\in\mathbb{Z}}(b_{4i}(X)-b_{4i+2}(X)).$
Proposition 3.2 states that $I_{JR}$ localizes as a signed sum over the fixed point set in a similar way to the signature. This puts the earlier results of [23, Theorem 1], [12, Corollary 6] into a natural context and generalizes them. The proof of Proposition 3.2 combines the localization formula for the signature (see Theorem 2.3), the localization formula for the Betti numbers (see Theorem 2.2), and Poincaré duality, along with some mild arithmetic simplifications.
In the next part of the article we use Proposition 3.2 often in combination with other techniques to prove some results which probe the relation between the Kähler and symplectic categories in the presence of a Hamiltonian $S^{1}$-action. The main thread is a collection of results regarding complete intersections which we now describe in more detail.
Symplectic torus actions on complete intersections. Subsection 3.1 and Section 4 are dedicated to the following problem.
Question 1.1.
Let $(M,\omega)$ be a closed symplectic manifold diffeomorphic to a complete intersection. Under which conditions does $(M,\omega)$ admit a Hamiltonian $S^{1}$-action?
@@
To summarize the results relevant to Question 1.1, we set the following notation: Let $X_{m}(d_{1},\ldots,d_{k})$ denote a smooth complete intersection in $\mathbb{CP}^{k+m}$ with multidegree $(d_{1},\ldots,d_{k})$. Here $m$ and $k$ are the the complex dimension and codimension respectively. For the definition of complete intersection see Definition 2.13, for more background material about complete intersections see Subsection 2.4.
@@
Since complete intersections of dimension at least $4$ are simply connected, there is no distinction between Hamiltonian and symplectic $S^{1}$-actions. For the sake of brevity, we keep the understanding that a complete intersection $(M,\omega)$ is a closed symplectic manifold diffeomorphic to a complete intersection. For some results appearing later in the article, some more restrictions on the symplectic form are needed and these are stated explicitly.
Note that in dimension $2$, $S^{2}$ and $T^{2}$ have symplectic circle actions with isolated fixed points which are Hamiltonian and non-Hamiltonian respectively, and smooth orientable surfaces of genus at least $2$ do not admit non-trivial smooth actions of $S^{1}$, this explains the inclusion $X_{1}(1)\cong X_{1}(2)\cong S^{2},X_{1}(3)\cong X_{1}(2,2)\cong T^{2}$ in the results below, and the exclusion of all other complete intersection curves. With this understanding, we work with complete intersections of dimension at least four where we may assume by simply connectedness that a symplectic action is Hamiltonian.
@@
The major motivating result of this paper is a theorem of Benoist [1, Theorem 3.1], which is a complete characterization of smooth complete intersections with infinite holomorphic automorphism groups. A comprehensive smooth analogue in real dimension $6$ was given by Dessai and Wiemeler [9]. One of the main aims of this article is to obtain a symplectic analogue of this theorem, specifically Theorem 1.2, Corollary 1.3 and Theorem 1.5.
@@
The first main result is to remove the technical assumption on the fixed point set in our previous result on complete intersections [29, Theorem 1.1] in dimension 8.
@@
Suppose that $(M,\omega)$ is an $8$-dimensional complete intersection with a symplectic $S^{1}$-action. Then $M$ is diffeomorphic to one of $X_{4}(1),X_{4}(2)$ or $X_{4}(2,2)$.
In fact we will deduce this result from the more general Theorem 3.6 which states that a closed symplectic $8$-manifold $M$ having a Hamiltonian $S^{1}$-action and satisfying $b_{2}(M)=1$ has positive definite intersection form. The proof is given in Subsection 3.1. The first step of the proof is to note that the condition $I_{JR}(M)=0$ is equivalent to the intersection form being positive definite. Then, since $\dim(M)=8$ all the terms of the right hand side of Equation (1) of Proposition 3.2 are of the form $\pm I_{JR}(F)$ where $F$ is a $4$-dimensional fixed component. The proof is concluded by proving that $I_{JR}(F)=0$ for all $4$-dimensional fixed components. This is proven in a different way depending on whether $F$ is extremal or not, and whether $F$ is contained in an isotropy $6$-manifold. The most difficult case when $F$ is non-extremal and not contained in an isotropy $6$-manifold, the desired condition is proved using the Duistermaat-Heckman theorem and a result from [36].
Theorem 3.6 may be considered as a result regarding a certain class of monotone symplectic $8$-manifolds with a Hamiltonian $S^{1}$-action, see Subsection 3.1 for more precise details. There has been a lot of recent progress towards classifying monotone symplectic manifolds with Hamiltonian symmetries, see [5, 3, 17, 30, 38] and the references contained therein.
In Section 4, we prove a classification result for symplectic $S^{1}$-actions on complete intersections under a condition on the fixed point set analogous to [29, Theorem 1.1], therefore removing the assumption on the dimension in op. cit.
@@
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic $S^{1}$-action with no fixed component $F$ satisfying $\dim(F)>0$ and $\dim(M)=\dim(F)\mod 4$. Then $M$ is diffeomorphic to one of $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ for some $m\in\mathbb{N}$ or $X_{1}(3)$, $X_{2}(3)$.
The proof of Theorem 1.3 for $n$ even is similar to [29, Theorem 1.1], applying Proposition 3.2 to obtain $I_{JR}(M)=0$ and then applying the classification of Libgober and Wood [28, Corollary 6.1]. For $n$ odd the proof is different, relying on the localization of Betti numbers formula for Hamiltonian $S^{1}$-actions (Theorem 2.2), known properties about the Betti numbers of complete intersections (Lemma 2.14), then finally applying the classification of complete intersections having vanishing middle Betti number due to Ewing and Moolgavkar [11, Corollary 5].
Theorem 1.3 has the following consequence.
@@
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic $S^{1}$-action having only isolated fixed points. Then $M$ is diffeomorphic to one of $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ for some $m\in\mathbb{N}$ or $X_{1}(3)$, $X_{2}(3)$.
Projective spaces and quadrics of any dimension admit a Hamiltonian torus action with only isolated fixed points. Each of the smooth $4$-manifolds underlying $X_{2}(2,2)$ and $X_{2}(3),$ admits a symplectic form invariant under a torus action, since either is diffeomorphic to $\mathbb{CP}^{2}$ blown-up in finitely many points. Hence, the only unknown case remaining from Theorem 1.4 is $X_{2m}(2,2)$ for $m\geq 2$.
For a complete intersection $M$ let $\omega_{FS}$ denote the restriction of the Fubini-Study form to $M$. In Subsection 4.2, we are able to deal with $X_{2m}(2,2)$ when the action is GKM, and the symplectic form is homotopic to $\omega_{FS}$. Thus obtaining in Theorem 1.5 a sharp classification of the possible multidegrees of complete intersections admitting such actions, as in Benoist’s original result on automorphism groups.
@@
Supposing the existence of such an action for a contradiction on $X_{2n}(2,2)$, the proof uses [14, Corollary 3.1] and [28, Proposition 3.4] to obtain an expression for $c_{1}c_{n-1}$ in terms of $n$ and [17, Lemma 4.3], to evaluate the integral of $c_{1}$ along a hypothetical toric $1$-skeleton Poincaré dual to $c_{n-1}$. This allows us to obtain an inequality contradicting the aforementioned expression for $c_{1}c_{n-1}$.
Unimodality of Betti numbers of closed symplectic manifolds admitting a Hamiltonian $S^{1}$-action. In the final section of the paper, we prove some new results regarding the unimodality of Betti numbers of a closed symplectic manifold admitting a Hamiltonian $S^{1}$-action. Tolman has asked whether for a closed symplectic manifold with a Hamiltonian $S^{1}$-action, the even Betti numbers are unimodal (i.e. weakly increasing up to the middle degree) [22, Problem 4.3]. For Kähler manifolds the property follows from the hard Lefschetz theorem. In [6] Cho and Kim gave an affirmative answer to [22, Problem 4.3 (1)] when $M$ is $8$-dimensional and the action has isolated fixed points. In Theorem 1.6 we prove the unimodality of even Betti numbers for $8$-dimensional manifolds having a Hamiltonian $S^{1}$-action with no extremal $4$-dimensional fixed components.
@@
Suppose that $(M,\omega)$ is a closed symplectic $8$-manifold having a Hamiltonian $S^{1}$-action, and neither $M_{\min}$ nor $M_{\max}$ is $4$-dimensional. Then $b_{2}(M)\leq b_{4}(M)$.
In Proposition 5.2 we prove that for a closed symplectic $12$-manifold with $b_{2}(M)=1$, satisfying a similar assumption on the fixed point set, the even Betti numbers are unimodal. To the author‘s knowledge this is new even under the assumption of isolated fixed points. Furthermore, in Proposition 5.3 we give weaker versions of Proposition 5.2 in general dimensions, generalizing an inequality of Cho [4] to more general fixed point sets. Finally, in Proposition 5.5 we combine this result with the Rokhlin-Ochanine theorem to obtain a final application.
Acknowledgments. I would like to thank Leonor Godinho and Silvia Sabatini for helpful discussions, in particular explaining to me the results of [17], which allowed me to deal with the intersection of two quadrics in the GKM case. I presented some of the results about unimodality of Betti numbers in the workshop “Perspectives in equivariant topology” in Cologne September 2022, I would like the thank those present for interesting discussions, in particular Liat Kessler and Susan Tolman. I would like to thank Dmitri Panov for collaborations and discussions which influenced the overall approach of the project and for comments on a draft version. I would like thank the referee for the detailed report which greatly improved the article.
@@ 2 Preliminaries
2.1 Hamiltonian $S^{1}$-actions
Let $\mathfrak{t}$ denote the Lie algebra of the compact torus $T$, and $\mathfrak{t}^{*}$ its dual. Recall that if a manifold $M$ has a smooth $T$-action, then for every $t\in\mathfrak{t}$ we can naturally associate an invariant vector field on $M$; $X_{t}$, tangent to the orbits of the action of the one-parameter family associated to $t$.
@@ 2.1 Hamiltonian $S^{1}$-actions
Let $(M,\omega)$ be a symplectic manifold with a smooth $T$-action by symplectomorphisms. The action is called Hamiltonian if there is a smooth $T$-invariant function $\mu:M\rightarrow\mathfrak{t}^{*}$ called the moment map such that
@@
The fixed point set of the action, denoted $M^{T}$ is equal to the critical set of $\mu$, and is a collection of symplectic submanifolds [33, Theorem 5.47]. A connected component of $M^{T}$ is called a fixed component. Moreover, there always exists a $T$-invariant, compatible almost complex structure.
When $T$ is $1$-dimensional, the moment map is referred to as the Hamiltonian and denoted $H:M\rightarrow\mathbb{R}$. The Hamiltonian $H$ is a perfect Morse-Bott function ([26, Page 34], see Theorem 2.2 below for a full statement), and a perfect Morse function if and only if $M^{S^{1}}$ is finite.
Let $G\subset S^{1}$ be a subgroup. The subset $M^{G}=\{x\in M|\;g.p=p\;\forall g\in G\}$ is an $S^{1}$-invariant symplectic submanifold for which the $S^{1}$-action is Hamiltonian, with Hamiltonian $H|_{M^{G}}$ [33, Lemma 5.53]. This will be used in the proof of Theorem 1.2. Connected components of $M^{G}$ are called isotropy submanifolds.
Let $J$ denote a $T$-invariant, compatible almost complex structure. Note that at $p\in M^{S^{1}}$, $T_{p}M$ admits the structure of a complex $S^{1}$-representation. Therefore, we may decompose $T_{p}M$ into irreducible representations:
$z.(z_{1},\ldots,z_{n})=(z^{w_{1}}z_{1},\ldots,z^{w_{n}}z_{n}),$
where $w_{i}\in\mathbb{Z}$ for each $i$. The multiset of integers $\{w_{i}\}$ is an invariant of the representation and called the weights of the $S^{1}$-action at $p$.
For a fixed component $F$ denote by $\lambda_{F}$ half the Morse-Bott index of $H$ along $F$. It holds that $\lambda_{F}$ is also equal to the the number of strictly negative weights of the $S^{1}$-action along $F$ counted with multiplicity [39, Page 7]. We will need the following localization formula for the Betti numbers due to Kirwan [26, Page 34]. We note that in [26] the result appears as an auxiliary result to prove that the norm squared of the moment map for a general Hamiltonian Lie group action is perfect, see [39, Section 2] for a comprehensive treatment focusing only on material relevant to proving Theorem 2.2 for $S^{1}$-actions.
@@
[26, Page 34] Let $(M,\omega)$ be a closed symplectic manifold having a Hamiltonian $S^{1}$-action. Then for each $i$,
$b_{i}(M)=\sum_{F\subset M^{S^{1}}}b_{i-2\lambda_{F}}(F),$
where the sum runs over each of the fixed components.
We use $\sigma(M)$ to denote the signature of $M$, which is equal to the signature of the symmetric bilinear form
$H^{\frac{\dim(M)}{2}}(M,\mathbb{R})\times H^{\frac{\dim(M)}{2}}(M,\mathbb{R})\rightarrow\mathbb{R},$
that is
$\sigma(M)=b_{\frac{\dim(M)}{2}}^{+}(M)-b_{\frac{\dim(M)}{2}}^{-}(M),$
where $b_{\frac{\dim(M)}{2}}^{+}(M)$ (resp. $b_{\frac{\dim(M)}{2}}^{-}(M)$) denotes the maximal subspace on which the form is positive (resp. negative) definite. $\sigma(M)$ is defined to be equal to zero for manifolds of dimension not divisible by $4$ and is equal to $1$ when $M$ is a single point.
The following localization formula is a consequence of the equivariant version of the Atiyah-Singer index theorem. The formula we will use is derived in [21, Section 5.8].
@@
2.2 GKM actions
GKM actions introduced in [18] are a natural class of torus actions in various categories, which having associated (multi)graph, usually called the GKM graph encoding topological and geometric information about the torus action. The amenability of the GKM graph to combinatorial arguments make GKM actions an attractive class of torus actions to study. In [17] results about GKM actions in the setting Hamiltonian torus actions on closed symplectic manifolds were proved, which we will use in Subsection 4.2. Firstly, we recall the definition.
Definition 2.4.
Let $(M,\omega)$ be a closed symplectic manifold with a Hamiltonian $T^{k}$-action. The action is called GKM if for every codimension $1$ subgroup $K\subset T^{k}$, $M^{K}$ has dimension at most $2$.
If such manifold has a GKM action then the subsets $M^{K}$ for each codimension $1$ subgroup $k$ form a collection of $\frac{n\chi(M)}{2}$, invariant symplectic $2$-spheres, corresponding to the edges of the GKM graph, whose union is called the toric 1 skeleton. To prove Theorem 4.2 we will need [17, Lemma 4.3], we restate it here for convenience.
Lemma 2.5.
[17, Lemma 4.3] Let $(M,\omega)$ be a closed symplectic $2n$-manifold with a Hamiltonian GKM $T^{k}$-action. Then the toric $1$-skeleton is Poincaré dual to $c_{n-1}$.
2.3 Applications of Duistermaat-Heckman
We begin by recalling the Duistermaat-Heckman theorem [10, Theorem 1.1]. Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action, and let $c$ be a regular value of $G$. Recall that
$M_{c}:=H^{-1}(c)/S^{1}$
has the structure of a symplectic orbifold, induced via projection from the coisotropic submanifold $H^{-1}(c)\subset M$, see [33, Page 175].
Before stating the Duistermaat-Heckman theorem, we have to recall some of the properties of the gradient flow for a Hamiltonian $S^{1}$-action. Fix an $S^{1}$-invariant compatible almost complex structure $J$, and let $g$ be the associated metric. Recall gradient vector field $\nabla_{g}H$, is the vector field $g$-dual to the $1$-form dHdH. Then, if $(a,b)$ is an interval on which $H$ is regular and $c,d\in(a,b)$ the flow of $Y$ gives an $S^{1}$-equivariant diffeomorphism $H^{-1}(a)\rightarrow H^{-1}(b)$ and therefore a orbifold diffeomorphism $F_{a,b}:M_{a}\rightarrow M_{b}$. The Duistermaat-Heckman theorem [10, Theorem 1.1] states that the variation of the cohomology
$F_{a,b}^{*}[\omega_{b}]\in H^{2}(M_{a},\mathbb{R})$
where $b$ varies and $a$ is constant is linear and gives an explicit formula for the direction of this linear path. As is usual in the subject, in the statement and applications of the Duistermaat-Heckman theorem that follow, we keep the above understanding of what variation of $\omega_{c}]$ means, and do not make explicit reference to the gradient flow.
Theorem 2.6.
[10, Theorem 1.1] Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action, with Hamiltonian $H$. Let $c$ be a regular value of $H$. Then
$\frac{d}{dt}[\omega_{t}]|_{t=c}=e(H^{-1}(c)).$
Here $H^{-1}(c)$ is considered as an (orbifold) principal $S^{1}$-bundle over the reduced symplectic orbifold $M_{c}$. Finally, $e$ denotes the Euler class of this bundle in $H^{2}(M_{t},\mathbb{R})$.
It is known that when $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action, all level sets of the Hamiltonian $H$ are connected [33, Lemma 5.51]. We denote by $M_{\min}$ and $M_{\max}$ the subsets on which the Hamiltonian obtains its minimum and maximum respectively. Note that since $H$ is critical along these subsets, they are pointwise fixed by $S^{1}$. Moreover, they are both connected, symplectic submanifolds by [33, Lemma 5.51] and [33, Theorem 5.47] respectively. The following lemma was proven in [39] using the Duistermaat-Heckman Theorem.
Lemma 2.7.
[39, Lemma 3.1] Let $(M,\omega)$ be closed symplectic manifold with a Hamiltonian $S^{1}$-action. Then for a fixed component $F$,
$\lambda_{F}\leq\sum_{H(F^{\prime})<H(F)}(\frac{\dim(F^{\prime})}{2}+1).$
We call a fixed point component $Y$ internal if $H(M_{\min})<H(Y)<H(M_{\max})$. Lemma 2.7 has the following consequence which we will use.
Lemma 2.8.
Suppose that $(M,\omega)$ is a closed symplectic manifold with a Hamiltonian $S^{1}$-action. Suppose that $M_{\min}$ is an isolated fixed. If $Y$ is an internal fixed point component such that $H(Y^{\prime})\geq H(Y)$ for any other internal fixed component $Y^{\prime}$, then $\lambda_{Y}=1$. If $Y$ is an internal fixed point component such that $H(Y^{\prime})\leq H(Y)$ for any other internal fixed component $Y^{\prime}$, then $\lambda_{Y}=n-1$.